A man goes 10 M towards north then 20m towards east then displacement is

Grade:11

1 Answers

Mahima Kanawat
5 years ago
Dear student 
10 m north and 20 m east (draw it), then displacement can be find out by Pythagoras i. e.,
D = ( 10+ 20)1/2 = 5001/2 = 10 (5)1/2 = 10*2.23 = 22.3 m
WITH REGARDS
MAHIMA 
ASKIITIANS FORUM EXPERT